OpenStudy (anonymous):

How are the sin 30° and the cos 60° related?

OpenStudy (anonymous):

the sine and cosine are (believe it or not) called "co-functions." One property of these "co-functions" is that sin A = cos (90 degrees - A). Note that 30 and 60 add up to 90. We call 30 and 60 "complementary angles," by virtue of their adding up to 90. If two angles add to 180, they are "supplementary angles," in case you were dying to know more than you have to to answer this particular homework problem. :) Therefore, sin 30 = cos (90-30) = cos 60.

OpenStudy (anonymous):

thank you so much